OpenStudy (anonymous):

ind the area of the region enclosed by these graphs and the vertical lines x = -1 and x = 2 f(x)=x+2 g(x)=x^2-4

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Interesting question. Could you show us what you've done so far please?

OpenStudy (anonymous):

i have to figure out which one is the upper and which one is the lower but I am not entire sure but I did find the graph

OpenStudy (anonymous):

Okay first you have to combine the two equations together. f(x)=x+2 g(x)=x^2-4 \[x+2=x^2-4\] \[x^2-x-4-2=0\] And you would get this right? \[x^2-x-6=0\]

OpenStudy (anonymous):

yes that is what I got

OpenStudy (anonymous):

Then you would put your end points. \[\int\limits_{-1}^{2}(x^2-x-6)dx\]

OpenStudy (anonymous):

Basically if you end up with a negative, just put absolute value signs. It won't matter because the negative number comes up due to the combinging process at the very beginning.

OpenStudy (anonymous):

So could you integrate that for me please?

OpenStudy (anonymous):

just a moment, workign it

OpenStudy (anonymous):

@RoseRoque Have you gotten an answer?

OpenStudy (anonymous):

hang on

OpenStudy (anonymous):

Is something wrong? Because this integral only requires anti-differentiation. Are you anti-differentiating or using another method because it's much easier to anti-differentiate integrals such as these than using other methods.

OpenStudy (anonymous):

-33/2

OpenStudy (anonymous):

look it takes me a minute because I do it all piece by piece by piece okay

OpenStudy (anonymous):

Okay, well in actual fact, it actually took 13-14 minutes. And unfortunately the answer is not -33/2. You must of done a mistake somewhere in your working.

OpenStudy (anonymous):

What did you get when you integrated the integral?

OpenStudy (anonymous):

Did you get this? \[=\frac{1}{3}x^3-\frac{1}{2}x^2-6x\]

OpenStudy (anonymous):

And it's good to have a calculator with you if you're not that good at adding/subtracting fractions. Makes everything easier for you and you don't need to work it out piece by piece, etc.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, well you should of gotten this. \[\left|=(\frac{8}{3}-2-12)-(-\frac{1}{3}-\frac{1}{2}+6)\right|\] \[\left|=\frac{9}{3}-8-\frac{1}{2}\right|\] \[\left| =-\frac{11}{2} \right|\]

OpenStudy (anonymous):

And then what would be your final answer. the absolute value signs weren't meant toinclude the equal signs inside. Sorry about that, butthe main quesion I want to ask is what would the final answer be?

OpenStudy (anonymous):

-33/2

OpenStudy (anonymous):

to include*

OpenStudy (anonymous):

but the*

OpenStudy (anonymous):

Using what I have as my working, what would be the final answer?

OpenStudy (anonymous):

i got the final answer to be -33/2

OpenStudy (anonymous):

-33/2 is incorrect. Are you sure you didn't make a mistake somewhere, because I double checked mine and I couldn't find a mistake.

OpenStudy (anonymous):

is it not supposed to be negative?

OpenStudy (anonymous):

Ah wait. Yes that's correct!

OpenStudy (anonymous):

Sorry about that. And yes, you're meant to put absoulte value signs around yourlines of working out.

OpenStudy (anonymous):

It doesn't really matter whether it's positive or negative, because if it is negative which is true in this case, you jsut put absolute value signs.

OpenStudy (anonymous):

so it is -33/2

OpenStudy (anonymous):

No, the answer has to be positive. But what you're jsut doing is making the number positive by putting absolute value signs around your number and then on the next line you write the number without the minus sign in front.

OpenStudy (austinl):

Here is how I would approach it. The top line, is the actual line, y=x+2. The bottom is the curve, y=x^2-4. So I would then set up the integral like this. You take the integral from -1 to 2. And you set it up, top minus bottom. \[\int\limits_{-1}^{2}(x-2)-(x^2-4) dx\] Which, when solved is 9/2. Did this help?

OpenStudy (austinl):

I apologize, I made a typo. It is supposed to be x+2 in the integral.

OpenStudy (noelgreco):

The correct answer is 33/2